Systems Staff Meeting Notes 12/13/99

Systems Staff Meeting Notes
December 13, 1999

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying Tang
  1. Bobbie found a 1926 Mercury dime over the weekend.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Virus scan - keep updated! Looks like there might be a massive onslaught of virus attacks over the New Year - hackers perceive opportunity with all of the Y2K hoopla. There are some attacks that can wipe out the HD & motherboard. Others are macro-viruses that are polymorphic, capable of mutating themselves. Be very careful about opening attachments. May be wise to turn off the preview pane in Outlook, due to a virus that is activated when previewed.
  2. Orders - Y2K upgrade for the Novell CD server - will upgrade it to 3.12 some time between Christmas & New Years.
  3. Various software & hardware parts are on order.
  4. 60 Cows enroute.
  5. Paint Shop Pro version 6 has arrived on metering software - 10 copies available, anxiously sought by DLC.
  6. We're going out to get the bid on a TerraByte server (trillion bytes, 1000 gig) for DLC. It is a SAN (Storage Attached Network) - fiber channel to be attached to the Dell server in the tower. The first TerraByte will cost somewhere in the neighborhood of 25 grand. We'll probably hold off on getting the second TerraByte until the price comes down. This server will be used to make available electronic texts in the DLC. It will be RAID upgradeable as well as expandable. The Libraries are looking for outside funding for this expenditure - a donor would get his/her name on the server. Order should be placed by the end of the year, probably won't be received until the spring.
  7. Evaluation of chatroom software for librarians - Carol Turner has placed an order for it. The idea is that it would serve as an online reference service. The particular model that this has been drawn from is one at the University of North Texas. We may have to put it up independently for security purposes - restricting access to within the Library. The status of this order is that it is still undergoing evaluation.
  1. NAC meeting this afternoon - topics include increasing the structure of authority - stemming from the CIO department. They are drafting a plan, consulting with the University attorney. Another topic will be the University's open mail relay policy - the problem with service configured this way is that people on the outside will bounce off the open relay - making it impossible to trace the real source. UF has been blacklisted in some places due to this policy.
  2. Smathers network reconfiguration - stalled.
  3. MSL Lighting Project - Ashcraft & crew will be moving the computers downstairs to the MSL conference room today. The switch is already there, we managed to convince them to set up only 14 machines (they wanted 18). The electrical power could be an issue - there are only two circuits. John Ashcraft will be out of the office for two weeks, but Michael Howell will be available. A Xerox printer will also be moved, which will pull a lot of power. The HVAC issue has not been addressed by MSL. We will be adding a new switch to improve access - the POP on the first floor will now be distributed from the 1st floor up to the 2nd, which is much more rational than the current setup. That switch is still located here in Systems - we've been trying to do a software upgrade, but have not been successful. Regardless, the switch will be functional without it.
  4. NT5 replacement - has been stable for the past 5-6 days, it's ready to go - just waiting for an opportune time to bring the old one down. Service pack 6a crashed the server, so we've backed off - only using service pack 4.
Software Projects Status
  1. Wiring program is close to finished - Ying received feedback from the techs on the demo. She is working on a system log, the multi-user problem and some other enhancements. May be finished by next week.
  2. NERDC tape conversions - sill ongoing, running into difficulties. Has been pending for about 3 weeks, may call Jim Arnette to speed up the process.
  3. General distribution of the Trouble Ticket Viewer - released to a few super liaisons for beta testing. Limited availability - tried to secure it to liaisons only, but the security change ended up locking everyone out. The viewer is web based & lightning fast.
  4. Form to mail - pending the Suzy/Winston server (SOMBA/Web server running under LINUX).
DP Coordinator
  1. Debra checked with Peter McKay regarding status of the CRSP Link database in use at the Business School. He said that it contains all the information they want. Peter just needs to verify that it is no longer in beta testing (that they are getting updates). When that is the case, Peter will notify Andy Halula at Compustat and let him know that we no longer need the cartridge or reel tapes as part of our regular subscription. At that point, Debra will be able to return the old tape shipments. The Library will continue to keep the site license for the Compustat & CRSP datafiles, although CRSP Link will actually be paid for & managed by the Business School. It is available on the Business School Network - any patrons looking for CRSP or Compustat data will now be referred to Peter McKay who will direct them to the appropriate Business School network source.
  1. Library Holiday Party - Wednesday, December 15, 2:00-4:00 PM. Three members of the Systems staff have been nominated for awards.
  2. Bobbie out Dec 16 & 17.
  3. Systems Department Christmas Luncheon - Friday, Dec. 17 - noonish. More details later this week.
  4. Rich out December 20 - 29.
  5. Bobbie out December 22-24.
  6. Will out December 28-29.
  7. New Year Holiday - Friday, Dec. 31.
  8. Everett out Jan. 6-12.
  9. Everett out Jan. 28-31.

Back to the Systems Homepage

Last updated December 20, 1999
by Debra Harris